OpenStudy (anonymous):

Use substitution to solve the system of equations. y + 1 = x -2x + 3y = 2 a) (5, 4) b) (-6, 3/2) c) (1/2, 15) d) (1/4, 1)

OpenStudy (marigirl):

your first equation is x=1+y what we can do is now 'substitute' 1+y instead of the x in the second equation. -2(1+y)+3y=2 expand the bracket -2-2y+3y=2 add two to both sides -2y+3y=4 y=4 we know now y=4 the first equation x=1+y so x=1+4 x=5 (5,4)